OpenStudy (pottersheep):

Integral of arctan(1/x). The (1/x) confuses me. I used integration by parts u = arctan(1/x), du= 1/(1+(1/x^2)) v=x, dv = 1 I see that my du is wrong, because of reverse chain rule or something, that's the part I don't get. Can someone explain to me what du would be please? Thank you

OpenStudy (aum):

Not reverse chain rule. Just straight chain rule. du= 1/(1+(1/x^2)) * d/dx (1/x)

myininaya (myininaya):

\[\text{ let } u=\arctan(\frac{1}{x}) \\ \text{ then } \tan(u)=\frac{1}{x}\\ \text{ so } sec^2(u) du=\frac{-1}{x^2} dx \\ -x^2 sec^2(u) du=dx\\\] \[-\cot^2(u)\sec^2(u)du=dx \\ -\csc^2(u) du =dx \] \[\int\limits_{}^{}\arctan(\frac{1}{x})dx=\int\limits_{}^{}-u \csc^2(u)du\] is probably the way I would go about it and you know then the integration by parts happens after this

myininaya (myininaya):

but you can go about it the way you have suggested

OpenStudy (pottersheep):

@aum gah I feel so silly! I was just complicated it for myself. Thank you both